Demand for a six-foot trees at Home Depot is approximated by a random variable with density function given below (units in hundred trees). Suppose Home Depot buys six-foot trees for $20 and sells them for $45. Any trees not sold are shredded. Find the number of trees that should be ordered to maximize expected profit.

cx, for 5 <= x <= 10

F(x)=

0, elsewhere
Answer #1

Solution

Back-up Theory

Expected profit = [Σ{Profit (p)} x P(p)], sum over all possible values of p.

Now, to work out the solution,

Let q = number of trees ordered and x = demand for trees.

Then, using the given density function, the numerical probabilities are as follows:

X (in hundreds)

5

6

7

8

9

10

Total

Probability density

5c

6c

7c

8c

9c

10c

45c

Numerical Probability

0.1111

0.1333

0.1556

0.1778

0.2000

0.2222

1.0000

Note that the total probability must be 1. Hence, 45c = 1 or c = 1/45.

Now, profit

= (45 - 20) = 25q if it sells, i.e., x ≥ q

= 25x - (q - x)(20) = 5x – 20q if x < q

Expected profit

q

x

5

6

7

8

9

10

5

125

125

125

125

125

125

6

105

150

150

150

150

150

7

85

130

175

175

175

175

8

65

110

155

200

200

200

9

45

90

135

180

225

225

10.00

25

70

115

160

205

250

Probability

0.1111

0.1333

0.1556

0.1778

0.2000

0.2222

Expd Profit

profit x probability - q = 5

13.889

16.667

19.444

22.225

25.000

27.778

125.003

6

11.667

20.000

23.333

26.670

30.000

33.333

145.003

7

9.444

17.333

27.222

31.115

35.000

38.889

159.004

8

7.222

14.667

24.111

35.560

40.000

44.444

166.004

9

5.000

12.000

21.000

32.004

45.000

50.000

165.004

10

2.778

9.333

17.889

28.448

41.000

55.556

155.004

Since expected profit of $16600.40 is maximum, optimum number of trees to be ordered is 8. Answer

DONE